The perimeter of a rectangle is 420 inches. The width is 4 inches less than the length. What is the length of the rectangle?
Answer by checkley75(3666)

W=L-4
2W+2L=420
2(L-4)+2L=420
2L-8+2L=420
4L=420+8
4L=428
L=428/4
L=107 IN. ANS. FOR THE LENGTH.
W=107-103=58 IN. ANS. FOR THE WIDTH.
PROOF:
2*107+2*103=240
214+206=420
420=420